OpenStudy (anonymous):

how do i solve 4^(3x)=12

OpenStudy (saifoo.khan):

\[\huge 3x = \log_4 12\]

OpenStudy (anonymous):

so how do i solve, Is that the answer

OpenStudy (saifoo.khan):

No. that is not the answer. use a calc to solve the right part.

OpenStudy (anonymous):

oh, K.

OpenStudy (anonymous):

@luisrock2008 Can you continue?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

Plug ln4/ ln 12 into your calculator, then divide the result for 3 to get x value.

OpenStudy (anonymous):

You mean ln 12 / ln 4

OpenStudy (anonymous):

@nbouscal, thanks, I'm about to fix it :)

OpenStudy (anonymous):

why ln

OpenStudy (anonymous):

That's how you change the base!

OpenStudy (anonymous):

\[ \begin{align} &\text{If}&y&=\log_ax,\\ &\text{then}&x&=a^y=e^{y\log a},\\ &\text{so}&\log x &=y\log a,\\ &\text{or}&y&=\frac{\log x}{\log a}.\\ &\text{In other words,}\\ &&\log_a x&=\frac{\log x}{\log a}. \end{align} \]

OpenStudy (anonymous):

all this is confusing, is there an equation I can use on any problems like this

OpenStudy (anonymous):

That is the equation you can use, but you want to learn why the equation works. Revisit the definitions of the exponential and logarithmic functions if necessary.

OpenStudy (anonymous):

Is there any website or youtube videos that can explain this all to me step by step.

OpenStudy (anonymous):

I'm sure Khan has some good videos on it, khanacademy.com